OpenStudy (anonymous):

square root of 500n over 4y^2

OpenStudy (anonymous):

can some one help

OpenStudy (mertsj):

\[\sqrt{\frac{500n}{4y^2}}\]

OpenStudy (mertsj):

Is that the problem?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is the answer 10 square root 5n over 2y

OpenStudy (mertsj):

\[\frac{\sqrt{500n}}{\sqrt{4y^2}}=\frac{\sqrt{100(5)n}}{2y}=\frac{10\sqrt{5n}}{2y}\]

OpenStudy (anonymous):

thank you so very much